# Mixed precision sketching for least-squares problems and its application in GMRES-based iterative refinement

### v1

@@ 1 Introduction
The goal of this work is twofold: to analyze the randomized sketching in two precisions and to determine when this approach to LSIR is guaranteed to converge, and give performance benefits. The theoretical analysis uses results for FGMRES backward stability and requires bounds for the condition number of the preconditioned augmented matrix.
@@ 2 Background on sketching
An embedding that has been theoretically analyzed and widely used is a Gaussian matrix, namely, $\Omega\in\mathbb{R}^{s\times m}$ with independent entries drawn at random from $\mathcal{N}(0,1/s)$ with an appropriately set $s$. Note that $\Omega$ is dense and thus for large problems storage and computing matrix-vector products can be expensive. It is known that $\Omega$ is an oblivious $(\epsilon,\delta,n)$-subspace embedding if $s=\Omega(\epsilon^{-2}\textrm{log}(n)\textrm{log}(1/\delta))$, however in some applications good results can be obtained by setting $s=n+5$; see, for example, [23, 27].
CountSketch is a sparse subspace embedding constructed by randomly and uniformly choosing one entry in each column of $\Omega$ and setting it to $1$ or $-1$ with probability $1/2$; all other entries in $\Omega$ are set to zero [12]. Such $\Omega$ thus randomly samples and adds/subtracts some rows of $A$. To ensure that $\Omega$ is an oblivious $(\epsilon,\delta,n)$-subspace embedding, $\Omega$ needs to have at least $(n^{2}+n)/\delta(2\epsilon-\epsilon^{2})^{2}$ rows. CountSketch thus requires more samples than a Gaussian embedding, although the latter is more expensive to apply. Sparse and dense embeddings can be combined to leverage this in what is called multisketching by applying two sketches in sequence; see, for example, [21, 28, 17] .
@@
where $A=U\Sigma V^{T}$ is the economic SVD and $A=Q_{A}R_{A}$ is the economic QR factorization; see [26, Theorem 1] and [24, Lemma 2.1]. Setting $P$ to be the $R$-factor of the economic QR decomposition $\Omega A=QR$ and using the economic QR decomposition $A=Q_{A}R_{A}$ we can also show that
$\|AR^{-1}\|_{2}=\|R_{A}R^{-1}\|_{2}=\|(\Omega Q_{A})^{\dagger}\Omega Q_{A}R_{A}R^{-1}\|_{2}\\ =\|(\Omega Q_{A})^{\dagger}QRR^{-1}\|_{2}\leq\|(\Omega Q_{A})^{\dagger}\|_{2},$
where $(\Omega Q_{A})^{\dagger}$ is the Moore-Penrose pseudoinverse of $\Omega Q_{A}$. The reduction of the norm and the condition number is thus determined by how well the sketching operator approximates the basis for the range of $A$. These quantities can be expressed via the subspace embedding distortion parameter $\epsilon$ as [22, Proposition 5.4]
@@ 3 Sketching in two precisions
where $c$ is a small constant that does not depend on $n$ [18, Section 2.2]. We make the standard assumption that no overflow or underflow occurs. In the following, hats denote computed quantities, that is, $\widehat{R}$ is the computed version of $R$. The notation $\lesssim$ is used when dropping terms that are negligible compared to other terms in the expression.
@@ 3.1 Computing $R$
$(2log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y})<1\quad\text{ and }$ (3)
$(2log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}<1$ (4)
@@
$\beta=$
$\left(1+(2log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\Omega A)+(2log_{2}n+4)^{2}n\widetilde{\gamma}_{sn}^{(QR)}\|\Delta_{s}\|_{2}\|(\Omega A)^{\dagger}\|_{2}\kappa_{2}(\Omega A)\right)$
×$\,\times\left(1+(2log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}\right),$
@@
$\|\widehat{R}\|_{2}$ $\lesssim\beta\|\Omega A\|_{2},$ (5)
$\|\widehat{R}^{-1}\|_{2}$ $\lesssim\beta\|(\Omega A)^{\dagger}\|_{2},$ (6)
$\kappa_{2}(\widehat{R})$ $\lesssim\beta^{2}\kappa_{2}(\Omega A),$ (7)
$\|A\widehat{R}^{-1}\|_{2}$ $\lesssim\beta\|AR^{-1}\|_{2},$ (8)
$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\lesssim\beta\|(AR^{-1})^{\dagger}\|_{2},$ (9)
$\kappa_{2}(A\widehat{R}^{-1})$ $\lesssim\beta^{2}\kappa_{2}(AR^{-1}).$ (10)
@@
We start with $\|\Gamma_{1}\|_{2}$, which requires considering the finite precision error in computing the QR decomposition. Standard results [18, Theorem 19.4] show that the Householder QR decomposition of $\widehat{Y}$ returns $\widehat{R}$ such that
@@
In [14, Theorem 3.1], it is shown that
$\|\Gamma_{1}\|_{2}\leq(2log_{2}n+4)\|R_{Y}^{-T}E_{c}R_{Y}^{-1}\|_{2}.$
@@
$\|\Gamma_{1}\|_{2}\leq(2log_{2}n+4)\left(n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y})+n(\widetilde{\gamma}_{sn}^{(QR)})^{2}\kappa_{2}(\widehat{Y})^{2}\right).$ (14)
@@
$\|\Gamma_{2}\|_{2}\leq$ $\,(2log_{2}n+4)\left(n^{1/2}\|R^{-1}\|_{2}\|\Delta_{s}\|_{2}+n\|R^{-1}\|_{2}^{2}\|\Delta_{s}\|_{2}^{2}\right)$
$=$ $\,(2log_{2}n+4)\left(n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}+n\|(\Omega A)^{\dagger}\|_{2}^{2}\|\Delta_{s}\|_{2}^{2}\right).$ (15)
@@
Taking the norms, using
$\|(I+\Gamma_{1})(I+\Gamma_{2})\|_{2}\leq(1+\|\Gamma_{1}\|_{2})(1+\|\Gamma_{2}\|_{2}),$
$\|(I-\Gamma_{2})(I-\Gamma_{1})\|_{2}\leq(1+\|\Gamma_{1}\|_{2})(1+\|\Gamma_{2}\|_{2}),$
ignoring the second order terms in (14) and (15), substituting
$\kappa_{2}(\widehat{Y})=\kappa_{2}(R_{Y})\leq\kappa_{2}(R)(1+\|\Gamma_{2}\|_{2})^{2}$
in (14), and
@@ 3.2 Comments
$(2log_{2}n+4)n^{1/2}m^{1/2}\gamma_{m+1}^{(s)}\frac{\|A\|_{2}\|\Omega\|_{2}}{\|\Omega A\|_{2}}\kappa_{2}(\Omega A)<1.$
@@
$\beta\lesssim 1+(2log_{2}n+4)n^{1/2}\left(\widetilde{\gamma}_{sn}^{(QR)}+m^{1/2}\gamma_{m+1}^{(s)}\frac{\|A\|_{2}\|\Omega\|_{2}}{\|\Omega A\|_{2}}\right)\kappa_{2}(\Omega A),$
where we ignore the $\mathcal{O}(u_{QR}u_{s})$ terms. A mixed precision implementation of Algorithm 3.1 in [16] uses $u_{QR}\leq u_{s}$, that is, the QR decomposition is computed in precision higher than in the sketching step. Note that our results do not suggest a higher quality approximation in this setting.
@@ 3.4 Sketching ill-conditioned matrices in low precision
$\|\widehat{R}\|_{2}$ $\lesssim\beta_{s}\|\Omega A_{s}\|_{2},$
$\|\widehat{R}^{-1}\|_{2}$ $\lesssim\beta_{s}\|(\Omega A_{s})^{\dagger}\|_{2},$
$\kappa_{2}(\widehat{R})$ $\lesssim\beta_{s}^{2}\kappa_{2}(\Omega A_{s}),$
$\|A\widehat{R}^{-1}\|_{2}$ $\lesssim\beta_{s}\|AR_{s}^{-1}\|_{2},$
$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\lesssim\beta_{s}\|(AR_{s}^{-1})^{\dagger}\|_{2},$
$\kappa_{2}(A\widehat{R}^{-1})$ $\lesssim\beta_{s}^{2}\kappa_{2}(AR_{s}^{-1}).$
@@
and $\|(AR_{s}^{-1})^{\dagger}\|_{2}$ as
@@ 4.2 IR convergence guarantees
$\frac{\|\widehat{w}_{i}-\widetilde{A}\delta\widehat{d}_{i}\|_{2}}{\|\widetilde{A}\|_{2}\|\delta\widehat{d}_{i}\|_{2}+\|\widehat{w}_{i}\|_{2}}\lesssim\frac{1.3c(n,k)}{1-\rho}\zeta\kappa_{2}(M_{L}),$
where $c(n,k)$ is a constant depending on $n$ and $k$, $\rho<1$ depends on $n$, the number of FGMRES iterations $k$, $u$, $u_{R}$, $M_{R}$, and $\widehat{Z}_{k}$ (the preconditioned basis matrix arising in FGMRES), and $\zeta$ depends on $u$, $u_{A}$, $u_{L}$, $M_{L}$, $M_{R}$, $\widetilde{A}$, $\widehat{Z}_{k}$, $\delta\widehat{d}_{i}$, and $\widehat{s}_{i}$; see [8] for further details.
@@
$\kappa_{2}(M_{L})\lesssim\beta\max\{\|\Omega A\|_{2},\|(\Omega A)^{\dagger}\|_{2},\beta\kappa_{2}(\Omega A)\}\eqqcolon\psi(\beta,\Omega A).$ (25)
@@ 5 Numerics for dense problems
The least-squares problem (1) is constructed with a synthetic dense $A$ generated as a ’randsvd’ matrix from the MATLAB test matrices gallery with $m=10^{3}$ and $n=10^{2}$, geometrically distributed singular values, and various choices of $\kappa_{2}(A)$. The right-hand side $b$ is a random vector with entries drawn from a uniform distribution in the interval $(0,1)$ and normalized to have a unit norm. Such a right-hand side gives $\|r\|_{2}=\|b-Ax\|_{2}\approx 1$ and thus the sensitivity of the least-squares problem depends on $\kappa_{2}(A)^{2}$; we note that when $\|r\|_{2}$ is small the sensitivity depends on $\kappa_{2}(A)$ instead; see, e.g., [18, Section 20.1]. We only test this large residual setting since it is the case where the augmented system approach to LSIR is expected to be most advantageous over other LSIR approaches in terms of refining the solution $x$; see, e.g., [7]. We generate the sketching matrix as $\Omega=(4n)^{-1/2}G$, where $G$ is a random $4n\times m$ matrix with entries drawn from a standard normal distribution. There is no scaling for the augmented system, that is, $\alpha=1$.
@@
We compute the condition numbers and norms of the preconditioned matrices. The results when QR in Algorithm 3.1 is computed in single precision (Tables 1 and 2) and in double precision (Tables 3, 4, and 5) show that as long as $\kappa_{2}(A)<u_{s}^{-1}$, preconditioning with $\widehat{R}$ keeps the condition numbers of both $M_{L}^{-1}\widetilde{A}M_{R}^{-1}$ and $A\widehat{R}^{-1}$ at $\mathcal{O}(1)$. If, however, $\kappa_{2}(A)>u_{s}^{-1}$, then $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ grows significantly due to the size of $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$; the value of $\|A\widehat{R}^{-1}\|_{2}$ always stays close to 2. These results agree with our theoretical observations. In the case when $\kappa_{2}(A)>u_{s}^{-1}$, $\widehat{R}$ is still effective in reducing $\kappa_{2}(A\widehat{R}^{-1})$ compared to $\kappa_{2}(A)$ by approximately a factor of $u_{s}^{-1}$. Note that as predicted by our analysis, we obtain $\widehat{R}$ of the same quality when we use both $u_{QR}=u_{s}$ and $u_{QR}<u_{s}$.
@@
The initial solves with LSQR give solutions and residuals of similar quality in the same or very similar number of iterations if $\kappa_{2}(A)$ is sufficiently smaller than $u_{s}^{-1}$; see Tables 6 and 7. If, however, this is not the case, we clearly obtain a worse preconditioner and LSQR needs significantly more iterations to converge or convergence is not reached in the preset number of iterations. Note that although an impractical number of iterations are required for $u_{s}$ set to a precision such that $\kappa_{2}(A)>u_{s}^{-1}$, the method is still able to reach the same accuracy as with $u_{s}$ set to a higher precision.
@@
LSIR convergence results with $u$ set to single and double are presented in Tables 8 and 9, respectively. Note that the iterative refinement process converges in the cases where our theoretical analysis holds, that is, when $\kappa_{2}(A)<u_{s}^{-1}$. In order to achieve convergence when $\kappa_{2}(A)$ is close to $u^{-1}$, we have to increase the precisions in FGMRES for computing the matrix-vector products and applying the preconditioner, and possibly allow more FGMRES iterations in every LSIR iteration; this is needed for FGMRES to reach the required backward error, see, e.g., [2] and [8]. We note that it is also possible to achieve LSIR convergence when $\kappa_{2}(A)>u_{s}^{-1}$ by significantly increasing the maximum number of FGMRES iterations, setting lower tolerance for FGMRES and/or increasing $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES. These combinations of $u_{s}$ and $\kappa_{2}(A)$ are however not covered by our analysis and the FGMRES parameters are highly problem dependent and require tuning in practice.
$\kappa_{2}(A)$ $\kappa_{2}(\widetilde{A})$ $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ $\|A\widehat{R}^{-1}\|_{2}$ $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\kappa_{2}(A\widehat{R}^{-1})$
1e+00 2.62e+00 7.44e+00 2.03e+00 1.46e+00 2.97e+00
1e+01 1.63e+02 7.44e+00 2.03e+00 1.46e+00 2.96e+00
1e+02 1.62e+04 7.45e+00 2.03e+00 1.46e+00 2.96e+00
1e+03 1.69e+06 9.29e+00 1.86e+00 1.74e+00 3.24e+00
1e+04 1.34e+08 3.40e+02 1.79e+00 1.19e+01 2.15e+01
1e+05 4.49e+08 2.48e+04 1.67e+00 1.05e+02 1.76e+02
1e+06 9.59e+08 1.80e+06 1.67e+00 8.97e+02 1.50e+03
1e+07 3.10e+08 1.57e+08 1.60e+00 8.48e+03 1.36e+04

### v2

@@ 1 Introduction
The goal of this work is twofold: to analyze the randomized sketching in two precisions and to determine when this approach to LSIR is guaranteed to converge, and give performance benefits. The theoretical analysis uses results for FGMRES backward stability and requires bounds for the condition number of the preconditioned augmented matrix. Our primary contributions are as follows.
  • We extend the finite precision analysis for generating a randomized preconditioner to the two precision case.
  • We provide the first theoretical proof that something other than a regular QR decomposition can provide a preconditioner of sufficient quality for LSIR to converge.
  • We provide the first theoretical proof that FGMRES with split preconditioning works for LSIR rather than GMRES restricted to left preconditioning.
@@ 2 Background on sketching
An embedding that has been theoretically analyzed and widely used is a Gaussian matrix, namely, $\Omega\in\mathbb{R}^{s\times m}$ with independent entries drawn at random from $\mathcal{N}(0,1/s)$ with an appropriately set $s$. Note that $\Omega$ is dense and thus for large problems storage and computing matrix-vector products can be expensive. Good quality results for sketching tall and skinny dense matrices are usually obtained by setting $s=cn$ with $c$ being small; see, for example, [17] where $c\leq 4$.
Another well-studied and frequently used embedding, which has a sparse structure, is a sparse sign matrix $\Omega=\sqrt{m/\zeta}\begin{pmatrix}\omega_{1}&\omega_{2}&\dots&\omega_{m}\end{pmatrix}$, where $\omega_{j}$ has $\zeta$ (with $2\leq\zeta\leq s$) entries chosen randomly and uniformly and set to $1$ or $-1$ with probability $1/2$; all other entries in $\omega_{j}$ are set to zero, see, e.g. [23, Section 9.2],[14],[12]. Such $\Omega$ thus randomly samples and adds/subtracts some rows of $A$. Although the theoretical results require more samples than for the Gaussian embeddings, reliable results are obtained with $s$ being a small multiple of $n$.
@@
where $A=U\Sigma V^{T}$ is the economic SVD and $A=Q_{A}R_{A}$ is the economic QR factorization; see [26, Theorem 1] and [24, Lemma 2.1]. We can also show that
$\|AP^{-1}\|_{2}=\|R_{A}P^{-1}\|_{2}=\|(\Omega Q_{A})^{\dagger}\Omega Q_{A}R_{A}P^{-1}\|_{2}\\ =\|(\Omega Q_{A})^{\dagger}QPP^{-1}\|_{2}=\|(\Omega Q_{A})^{\dagger}\|_{2},$
where $(\Omega Q_{A})^{\dagger}$ is the Moore-Penrose pseudoinverse of $\Omega Q_{A}$ and the last equality is due to the columns of $Q$ being the orthonormal basis for $\textrm{range}(\Omega Q_{A})$. The reduction of the norm and the condition number is thus determined by how well the sketching operator approximates the basis for the range of $A$. In our case, we set $P$ to be the $R$-factor of the economic QR decomposition $\Omega A=QR$. The norm reduction can be expressed via the subspace embedding distortion parameter $\epsilon$ as [22, Proposition 5.4]
@@ 3 Sketching in two precisions
where $c$ is a small constant that does not depend on $n$ [19, Section 2.2]. We make the standard assumption that no overflow or underflow occurs. In the following, hats denote computed quantities, that is, $\widehat{R}$ is the computed version of $R$.
@@ 3.1 Computing $R$
$(2\log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y})<1\quad\text{ and }$ (3)
$(2\log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}<1$ (4)
@@
$\beta=$ $\,\left(1+18(2\log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\Omega A)\right)$
×$\,\times\left(1+2(2\log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}\right),$
@@
$\|\widehat{R}\|_{2}$ $\leq\beta\|\Omega A\|_{2},$ (5)
$\|\widehat{R}^{-1}\|_{2}$ $\leq\beta\|(\Omega A)^{\dagger}\|_{2},$ (6)
$\kappa_{2}(\widehat{R})$ $\leq\beta^{2}\kappa_{2}(\Omega A),$ (7)
$\|A\widehat{R}^{-1}\|_{2}$ $\leq\beta\|AR^{-1}\|_{2},$ (8)
$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\leq\beta\|(AR^{-1})^{\dagger}\|_{2},$ (9)
$\kappa_{2}(A\widehat{R}^{-1})$ $\leq\beta^{2}\kappa_{2}(AR^{-1}).$ (10)
@@
We start with $\|\Gamma_{1}\|_{2}$, which requires considering the finite precision error in computing the QR decomposition. Standard results [19, Theorem 19.4] show that the Householder QR decomposition of $\widehat{Y}$ returns $\widehat{R}$ such that
@@
In [15, Theorem 3.1], it is shown that
$\|\Gamma_{1}\|_{2}\leq(2\log_{2}n+4)\|R_{Y}^{-T}E_{c}R_{Y}^{-1}\|_{2}.$
@@
$\|\Gamma_{1}\|_{2}\leq(2\log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y})\left(1+n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y})\right)\eqqcolon\alpha_{1}.$ (14)
@@
$\|\Gamma_{2}\|_{2}\leq$ $\,(2\log_{2}n+4)\left(n^{1/2}\|R^{-1}\|_{2}\|\Delta_{s}\|_{2}+n\|R^{-1}\|_{2}^{2}\|\Delta_{s}\|_{2}^{2}\right)$
$=$ $\,(2\log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}\left(1+n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2}\right)\eqqcolon\alpha_{2}.$ (15)
@@
Note that $(A\widehat{R}^{-1})^{\dagger}=\widehat{R}A^{\dagger}$, because $A$ is full rank and $\widehat{R}$ is invertible. Taking the norms, using
$\|(I+\Gamma_{1})(I+\Gamma_{2})\|_{2}\leq(1+\|\Gamma_{1}\|_{2})(1+\|\Gamma_{2}\|_{2})\leq(1+\alpha_{1})(1+\alpha_{2}),$
$\|(I-\Gamma_{2})(I-\Gamma_{1})\|_{2}\leq(1+\|\Gamma_{1}\|_{2})(1+\|\Gamma_{2}\|_{2})\leq(1+\alpha_{1})(1+\alpha_{2}),$
noting that under the assumptions (3) and (4)
$\alpha_{1}$ $\leq 2(2\log_{2}n+4)n^{1/2}\widetilde{\gamma}_{sn}^{(QR)}\kappa_{2}(\widehat{Y}),$
$\alpha_{2}$ $\leq 2(2\log_{2}n+4)n^{1/2}\|(\Omega A)^{\dagger}\|_{2}\|\Delta_{s}\|_{2},$
and substituting $\kappa_{2}(\widehat{Y})$ in the bound for $\alpha_{1}$ by
$\kappa_{2}(\widehat{Y})=\kappa_{2}(R_{Y})\leq\kappa_{2}(R)(1+\|\Gamma_{2}\|_{2})^{2}\leq\kappa_{2}(R)(1+\alpha_{2})^{2},$
where using (4) we have $(1+\alpha_{2})^{2}\leq 9$, and using
@@ 3.2 Comments
$(2\log_{2}n+4)n^{1/2}m^{1/2}\gamma_{m+1}^{(s)}\frac{\|A\|_{2}\|\Omega\|_{2}}{\|\Omega A\|_{2}}\kappa_{2}(\Omega A)<1.$
@@
$\beta\leq 1+2(2\log_{2}n+4)n^{1/2}\left(9\widetilde{\gamma}_{sn}^{(QR)}+m^{1/2}\gamma_{m+1}^{(s)}\frac{\|A\|_{2}\|\Omega\|_{2}}{\|\Omega A\|_{2}}\right)\kappa_{2}(\Omega A)+\mathcal{O}(u_{QR}u_{s}).$
A mixed precision implementation of Algorithm 3.1 in [17] uses $u_{QR}\leq u_{s}$, that is, the QR decomposition is computed in precision higher than in the sketching step. Note that our results do not suggest a higher quality approximation in this setting.
@@ 3.4 Sketching ill-conditioned matrices in low precision
$\|\widehat{R}\|_{2}$ $\leq\beta_{s}\|\Omega A_{s}\|_{2},$
$\|\widehat{R}^{-1}\|_{2}$ $\leq\beta_{s}\|(\Omega A_{s})^{\dagger}\|_{2},$
$\kappa_{2}(\widehat{R})$ $\leq\beta_{s}^{2}\kappa_{2}(\Omega A_{s}),$
$\|A\widehat{R}^{-1}\|_{2}$ $\leq\beta_{s}\|AR_{s}^{-1}\|_{2},$
$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$ $\leq\beta_{s}\|(AR_{s}^{-1})^{\dagger}\|_{2},$
$\kappa_{2}(A\widehat{R}^{-1})$ $\leq\beta_{s}^{2}\kappa_{2}(AR_{s}^{-1}).$
@@
and using the assumption that $A$ is full rank and thus has linearly independent columns while $R_{s}$ has linearly independent rows we bound $\|(AR_{s}^{-1})^{\dagger}\|_{2}$ as
@@ 4.2 IR convergence guarantees
$\frac{\|\widehat{w}_{i}-\widetilde{A}\delta\widehat{d}_{i}\|_{2}}{\|\widetilde{A}\|_{2}\|\delta\widehat{d}_{i}\|_{2}+\|\widehat{w}_{i}\|_{2}}\leq\frac{1.3c(n,k)}{1-\rho}\zeta\kappa_{2}(M_{L}),$
where the second order terms are ignored, $c(n,k)$ is a constant depending on $n$ and $k$, $\rho<1$ depends on $n$, the number of FGMRES iterations $k$, $u$, $u_{R}$, $M_{R}$, and $\widehat{Z}_{k}$ (the preconditioned basis matrix arising in FGMRES), and $\zeta$ depends on $u$, $u_{A}$, $u_{L}$, $M_{L}$, $M_{R}$, $\widetilde{A}$, $\widehat{Z}_{k}$, $\delta\widehat{d}_{i}$, and $\widehat{s}_{i}$; see [8] for further details. We note that $\zeta$ is expected to dominate the term $\frac{1.3c(n,k)}{1-\rho}\zeta$, and the achievable backward error depends on the interaction between the precisions $u$, $u_{A}$, and $u_{L}$, and the error in applying $M_{L}^{-1}$ and computing the matrix-vector products with $\widetilde{A}$; guidance on how to achieve $\frac{1.3c(n,k)}{1-\rho}\zeta=\mathcal{O}(u)$ is provided in [8, Section 2.2].
@@
$\kappa_{2}(M_{L})\leq\beta\max\{\|\Omega A\|_{2},\|(\Omega A)^{\dagger}\|_{2},\beta\kappa_{2}(\Omega A)\}\eqqcolon\psi(\beta,\Omega A).$ (25)
@@ 5 Numerics for dense problems
The least-squares problem (1) is constructed with a synthetic dense $A$ generated as a ’randsvd’ matrix from the MATLAB test matrices gallery with $m=10^{3}$ and $n=10^{2}$, geometrically distributed singular values, and various choices of $\kappa_{2}(A)$. The right-hand side $b$ is a random vector with entries drawn from a uniform distribution in the interval $(0,1)$ and normalized to have a unit norm. Such a right-hand side gives $\|r\|_{2}=\|b-Ax\|_{2}\approx 1$ and thus the sensitivity of the least-squares problem depends on $\kappa_{2}(A)^{2}$; we note that when $\|r\|_{2}$ is small the sensitivity depends on $\kappa_{2}(A)$ instead; see, e.g., [19, Section 20.1]. We only test this large residual setting since it is the case where the augmented system approach to LSIR is expected to be most advantageous over other LSIR approaches in terms of refining the solution $x$; see, e.g., [7]. We generate the sketching matrix as $\Omega=(4n)^{-1/2}G$, where $G$ is a random $4n\times m$ matrix with entries drawn from a standard normal distribution. There is no scaling for the augmented system, that is, $\alpha=1$.
@@
We compute the condition numbers and norms of the preconditioned matrices. The results when QR in Algorithm 3.1 is computed in single precision (Figure 3) and in double precision (Figure 4) show that as long as $\kappa_{2}(A)<u_{s}^{-1}$, preconditioning with $\widehat{R}$ keeps the condition numbers of both $M_{L}^{-1}\widetilde{A}M_{R}^{-1}$ and $A\widehat{R}^{-1}$ at $\mathcal{O}(1)$. If, however, $\kappa_{2}(A)>u_{s}^{-1}$, then $\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$ grows significantly due to the size of $\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$; the value of $\|A\widehat{R}^{-1}\|_{2}$ always stays close to 2. These results agree with our theoretical observations. In the case when $\kappa_{2}(A)>u_{s}^{-1}$, $\widehat{R}$ is still effective in reducing $\kappa_{2}(A\widehat{R}^{-1})$ compared to $\kappa_{2}(A)$ by approximately a factor of $u_{s}^{-1}$. Note that as predicted by our analysis, we obtain $\widehat{R}$ of the same quality when we use both $u_{QR}=u_{s}$ and $u_{QR}<u_{s}$.
@@
The initial solves with LSQR give solutions and residuals of similar quality in the same or very similar number of iterations if $\kappa_{2}(A)$ is sufficiently smaller than $u_{s}^{-1}$; see Tables 1 and 2. If, however, this is not the case, we clearly obtain a worse preconditioner and LSQR needs significantly more iterations to converge or convergence is not reached in the preset number of iterations. Note that although an impractical number of iterations are required for $u_{s}$ set to a precision such that $\kappa_{2}(A)>u_{s}^{-1}$, the method is still able to reach the same accuracy as with $u_{s}$ set to a higher precision.
@@
LSIR convergence results with $u$ set to single and double are presented in Tables 3 and 4, respectively. Note that the iterative refinement process converges in the cases where our theoretical analysis holds, that is, when $\kappa_{2}(A)<u_{s}^{-1}$. In order to achieve convergence when $\kappa_{2}(A)$ is close to $u^{-1}$, we have to increase the precisions in FGMRES for computing the matrix-vector products and applying the preconditioner, and possibly allow more FGMRES iterations in every LSIR iteration; this is needed for FGMRES to reach the required backward error, see, e.g., [2] and [8]. We note that it is also possible to achieve LSIR convergence when $\kappa_{2}(A)>u_{s}^{-1}$ by significantly increasing the maximum number of FGMRES iterations, setting lower tolerance for FGMRES and/or increasing $u_{A}$, $u_{L}$ and $u_{R}$ in FGMRES. These combinations of $u_{s}$ and $\kappa_{2}(A)$ are however not covered by our analysis and the FGMRES parameters are highly problem dependent and require tuning in practice.
$10^{0}$$10^{2}$$10^{4}$$10^{6}$$10^{0}$$10^{3}$$10^{6}$$10^{9}$$\kappa_{2}(A)$$\kappa_{2}(M_{L}^{-1}\widetilde{A}M_{R}^{-1})$$\|A\widehat{R}^{-1}\|_{2}$$\|(A\widehat{R}^{-1})^{\dagger}\|_{2}$$\kappa_{2}(A\widehat{R}^{-1})$$\kappa_{2}(\widetilde{A})$
(a) $u_{s}$ half, $u_{QR}$ single
[…]
(b) $u_{s}$ single, $u_{QR}$ single